# LLMs can Perform Multi-Dimensional Analytic Writing Assessments: A Case Study of L2 Graduate-Level Academic English Writing

### v1

@@ A.2 Details of the 5-Unit Tutorial Series
To support their writing, the authors were provided with a short, curated bibliography for each task, designed to help them focus on literature review writing while minimizing the effort required for bibliographic searches. Prior to submitting their final writing samples for expert assessments, the authors engaged in peer reviews (for topics 1, 3, and 5) or group collaboration (for topics 2 and 4).
@@ A.3 Assessment Criteria
4. Academic vocabulary (academic) formulaic expressions and their functions Interactive e-book, Discussion forum, quiz 2 weeks Collaborative writing on lessons from the COVID- 19 pandemic
@@ B.1 Implementation
Table 8: An example of implementing our framework on a feedback comment on the second assessment criteria, i.e., material integration & citation, which is about citation practices. Our framework correctly identifies three problems identified in the comment and performs Problem Classification on each one of them, of which the first problem offers a concrete correction. The Correction Relevance Check is thus only performed on the first extracted problem.
@@ Appendix C Results
C.1 Scores
@@
Second, we observe that human assessors assigned different scores to identical or similar comments, mostly within 1-point differences. For example, assessor F gave the same comment “Decent number of citations” three times but assigned three different scores: 6, 7, and 8. Similarly, assessor C assigned scores of 7 and 8 to the comment “Appropriate use of connectors.” However, when the same comment is repeated, scores tend to be very close, typically within one point. For instance, assessor A assigned a score of 8 to the comment “Great use of academic words and formal tone” five times, with only one instance where the score was 9.
@@ C.2 Comments
C.3 Score-Comment Interaction
@@ Appendix E Prompts
E.1 Prompts for the Feedback Comment Quality Evaluation Framework Pipeline

### v2

@@ Acknowledgments
Zhengxiang Wang and Owen Rambow were supported in part by funding from the Defense Advanced Research Projects Agency (DARPA) under Contracts No.HR01121C0186, No. HR001120C0037, and PR No. HR0011154158. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA.
@@
We thank Yongjun Zhang and the three anonymous reviewers for their valuable feedback. This work was presented at several venues, including All Things Language and Computation (ATLAC) at Stony Brook University, the Mid-Atlantic Student Colloquium on AI, Language, and Learning (MASC-ALL) at Penn State University, and the New England NLP Meeting Series (NENLP) at Yale University. We are grateful for the insightful discussions and feedback received from the audiences at these events.
@@ A.1 Basic Corpus Statistics
T1 T2 T3 T4 T5
# Essays 50 16 31 13 31
Avg WC (w/o refs) 845 1169 926 1079 887
Avg WC (w/ refs) 1232 1583 1347 1666 1159
Table 5: Basic statistics of the corpus. “T” in each column stands for “Topic.” “WC” means “word count.”
@@ A.3 Assessment Criteria
4. Academic vocabulary Academic formulaic expressions and their functions Interactive e-book, Discussion forum, quiz 2 weeks Collaborative writing on lessons from the COVID- 19 pandemic
@@ B.1 Implementation
Table 8: An example implementation of our framework ProEval on a feedback comment on the second assessment criteria, i.e., material integration & citation, which is about citation practices. ProEval correctly identifies three problems (color-coded) identified in the comment and performs Problem Classification on each one of them, of which only the first problem offers a concrete correction. The Correction Relevance Check is thus only performed on the first extracted problem.
@@ Appendix C Results
C.1 Number of Co-Assessed Essays
Table 11 shows the number of essays co-assessed by different assessor pairs.
Human B Human C Human F Llama-3 (IM 1) Gemini-1.5 (IM 1) Llama-3 (IM 2)
Human B 141 93 106 140 140 139
Human C 93 93 78 93 92 93
Human F 106 78 106 106 105 106
Llama-3 (IM 1) 140 93 106 140 139 139
Gemini-1.5 (IM 1) 140 92 105 139 140 138
Llama-3 (IM 2) 139 93 106 139 138 139
Table 11: Number of essays co-assessed by different assessor pairs. We only show three LLMs, which failed to generate assessments for all 141 essays in the corpus for some reason (e.g., content moderation, exceeding context window). We omit the other LLMs since they assessed all essays and the numbers of essays they co-assessed with the five assessors in the table excluding human B are identical to those between human B and those five assessors. The number of essays the omitted LLMs and human B co-assessed is always 141.
C.2 Scores
@@
Second, we observe that human assessors assigned different scores to identical or similar comments, mostly within 1-point differences. For example, assessor F gave the same comment “Decent number of citations” three times but assigned three different scores: 6, 7, and 8. Similarly, assessor C assigned scores of 7 and 8 to the comment “Appropriate use of connectors.” However, when the same comment is repeated, scores tend to be very close, typically within one point. For instance, assessor A assigned a score of 8 to the comment “Great use of academic words and formal tone” five times, and there was only one more instance where the score was 9.
@@ C.3 Comments
C.4 Score-Comment Interaction
@@ Appendix E Prompts
E.1 Prompts for the Feedback Comment Quality Evaluation Framework ProEval